OpenStudy (anonymous):

if p = sin40 then why is sin140 = p too? isnt it supposed to be a 90 degree difference or something?

OpenStudy (anonymous):

\[ \sin(\pi - x) = \sin(x)\]

OpenStudy (anonymous):



OpenStudy (anonymous):

Take x=40

OpenStudy (anonymous):

nope it is because first we took 0+40 degree the next time we take 180-40 degree in use.. ok

OpenStudy (anonymous):